OpenStudy (anonymous):

Which solution set best represents X minus 6 is greater than or equal to twenty-five? a solution set contains thirty-one, thirty-two, thirty-three, and so on a solution set contains thirty-three, thirty-two, thirty-one, and so on a solution set contains thirty-two, thirty-three, thirty-four, and so on a solution set contains thirty-two, thirty-one, thirty, and twenty-nine

OpenStudy (anonymous):

the result must be greater than or equal to 25. every input is decreased -6, so if the independent variable x is 25, the result is less

OpenStudy (debbieg):

you need: \(\Large x-6 \ge 25\) You solve an inequality just like you do an equality (isolate the x), with ONE additional rule: if you MULTIPLY or DIVIDE by a negative number, you switch the direction of the inequality. So apply that here and see what your solution is.

OpenStudy (debbieg):

However, I'm puzzled by your solution sets.... is x limited to being an integer??

OpenStudy (anonymous):

I coped the question right from the work so i guess x is the limited to being an integer.

OpenStudy (debbieg):

So it actually says: "a solution set contains thirty-one, thirty-two, thirty-three, and so on"? Or does it show you a solution set (for the first one, for example): \[\Large \left\{ 31, 32, 33... \right\}\] ?

OpenStudy (anonymous):

yeah it shows that

OpenStudy (debbieg):

Either way, there is clearly a correct (or at least "best") answer. :) Just not what I would have expected for the solution sets, had it not been multiple choice, lol.

OpenStudy (debbieg):

So what do you get for the solution to the inequality?

OpenStudy (debbieg):

\(\Large x-6 \ge 25\)

OpenStudy (anonymous):

The first one?

OpenStudy (debbieg):

Right. Just add 6 to both sides, you get \(\Large x \ge 31\)

OpenStudy (debbieg):

Which generally would mean ALL REAL NUMBERS 31 and greater, not just integers.... but since that isn't one of your choices, I would say the first solution set is the "most correct". Maybe somewhere for the problem set you were told that X represents an integer?

OpenStudy (anonymous):

May I answer too ?:) And join :)

OpenStudy (anonymous):

sure if you wanna

OpenStudy (anonymous):

Math is so Confusing -_-

OpenStudy (anonymous):

I cant think this :)

OpenStudy (debbieg):

I thought it was already answered? lol

OpenStudy (anonymous):

So the answer is the second one.......Right?

OpenStudy (anonymous):

Wait ....

OpenStudy (anonymous):

So im wrong?

OpenStudy (anonymous):

wait its the first one

OpenStudy (debbieg):

Yes, the first one. I said above that was right. :) Do you understand WHY?